# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 4. Koszul duality
In this section we study Koszul duality between $\infty$-operads and $\infty$-cooperads and their respective $\infty$-categories of algebras and coalgebras. We construct this Koszul duality from Koszul-duality between augmented associative algebras and coaugmented coassociative coalgebras and their respective $\infty$-categories of modules and comodules following [23, 5.2.], [3]. The next two theorems are [3, Theorem 3.26.]:
@@
[…]
@@
Square (10) induces on the fiber over every ${\mathrm{A}}\in\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}$ a functor
@@
${\mathrm{Bar}}_{\mathcal{C}}:\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\rightleftarrows{\mathrm{coAlg}}({\mathcal{C}})_{{\mathbb{1}}/}:{\mathrm{Cobar}}_{\mathcal{C}}:={\mathrm{Bar}}^{\mathrm{op}}_{{\mathcal{C}}^{\mathrm{op}}},$
${\mathrm{Bar}}_{\mathcal{D}}:{\mathrm{BMod}}({\mathcal{D}})\rightleftarrows\mathrm{coBMod}({\mathcal{D}}):{\mathrm{Cobar}}_{\mathcal{D}}:={\mathrm{Bar}}^{\mathrm{op}}_{{\mathcal{D}}^{\mathrm{op}}}$
and square (10) is a map of adjunctions and so induces for any augmented algebra ${\mathrm{A}}$ in ${\mathcal{C}}$ and map ${\mathrm{Bar}}_{\mathcal{C}}({\mathrm{A}})\to{\mathrm{A}}^{\prime}$ of coaugmented coalgebras in ${\mathcal{C}}$ an adjunction
@@
${\mathbb{1}}\otimes_{\mathrm{A}}\mathrm{triv}_{{\mathrm{A}}}({\mathrm{Y}})\simeq\mathrm{colim}_{{\mathrm{n}}\in\Delta^{\mathrm{op}}}({\mathrm{A}}^{\otimes{\mathrm{n}}}\otimes{\mathrm{Y}})\to{\mathrm{Bar}}_{\mathcal{C}}({\mathrm{A}})\otimes{\mathrm{Y}}\simeq\mathrm{colim}_{{\mathrm{n}}\in\Delta^{\mathrm{op}}}({\mathrm{A}}^{\otimes{\mathrm{n}}})\otimes{\mathrm{Y}}$
@@
    $(-)^{\vee}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{/{\mathrm{triv}}}\simeq\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})_{/{\mathrm{triv}}}\rightleftarrows{\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})_{{\mathrm{triv}}/}\simeq\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}_{{\mathrm{triv}}/}:(-)^{\vee}.$
@@
    There is an adjunction ${\mathrm{LMod}}({\mathcal{C}})\rightleftarrows{\mathrm{coLMod}}({\mathcal{C}})$ and a map of adjunctions
    ${{\mathrm{LMod}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{coLMod}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{/{\mathrm{triv}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}_{{\mathrm{triv}}/}.}$
@@
    $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})={\mathrm{LMod}}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})={\mathrm{coLMod}}_{{\mathcal{Q}}}({\mathcal{C}}),$
    where the left adjoint lifts the functor ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ left adjoint to the trivial ${\mathcal{O}}$-algebra functor and the right adjoint lifts the functor ${\mathrm{triv}}\circ^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the trivial ${\mathcal{Q}}$-coalgebra functor.
@@
    $(-)^{\vee}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{/{\mathrm{triv}}}\rightleftarrows\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}_{{\mathrm{triv}}/}:(-)^{\vee}.$
@@
    ${{\mathrm{RMod}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{coRMod}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{/{\mathrm{triv}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}_{{\mathrm{triv}}/}.}$
@@
    ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\rightleftarrows{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}),$
    where the left adjoint lifts the functor $(-)\circ_{\mathcal{O}}{\mathrm{triv}}:{\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\to{\mathcal{C}}$ left adjoint to the trivial right ${\mathcal{O}}$-module functor and the right adjoint lifts the functor $(-)\circ^{{\mathcal{Q}}}{\mathrm{triv}}:{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\to{\mathcal{C}}$ right adjoint to the trivial right ${\mathcal{Q}}$-comodule.
@@
    ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}}):{\mathrm{triv}}*^{{\mathcal{Q}}}(-),$
    where the left adjoint lifts the functor ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ left adjoint to the trivial ${\mathcal{O}}$-algebra functor and the right adjoint lifts the functor ${\mathrm{triv}}*^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the trivial ${\mathcal{Q}}$-coalgebra functor.
@@
${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}}):{\mathrm{triv}}\circ^{{\mathcal{Q}}}(-),$
where the left adjoint lifts the functor ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ left adjoint to the trivial ${\mathcal{O}}$-algebra functor and the right adjoint lifts the functor ${\mathrm{triv}}\circ^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the trivial ${\mathcal{Q}}$-coalgebra functor. The forgetful functor ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})$ admits a right adjoint ${\mathrm{R}}.$ The functor
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ_{\mathcal{O}}(-)}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})\xrightarrow{{\mathrm{R}}}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ^{\mathcal{Q}}(-)}\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}).$
The latter functor lifts the functor ${\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the composition
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}^{\mathrm{dp},{\mathrm{conil}}}_{\mathrm{Q}}}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}}),$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$
lifts the functor ${\mathrm{triv}}*^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to ${\mathrm{triv}}_{\mathcal{Q}}.$
@@
${\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})\subset{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}^{\prime})\subset{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}^{\prime\prime}).$
@@
${\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})\subset{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}^{\prime})$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}^{\prime\prime})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}^{\prime\prime})$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$
@@
${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}^{\prime\prime})\rightleftarrows{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}}^{\prime\prime})$
@@
${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}^{\prime})\rightleftarrows{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}}^{\prime})$
@@
$\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\subset\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}^{\prime})\xrightarrow{{\mathrm{triv}}\circ_{{\mathcal{O}}}(-)}{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}}^{\prime})\xrightarrow{{\mathrm{L}}^{\prime\prime\prime}}{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})$
is left adjoint to the functor ${\mathrm{triv}}\ast^{{\mathcal{Q}}}(-):{\mathrm{coAlg}}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}).$
@@
(11) $(-)\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{triv}}:{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\rightleftarrows\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}):(-)\circ^{{\mathcal{Q}}}{\mathrm{triv}}$

### v2

@@ 3. Koszul duality
In this section we study Koszul duality between $\infty$-operads and $\infty$-cooperads and their respective $\infty$-categories of algebras and coalgebras. We construct this Koszul duality from Koszul-duality between augmented associative algebras and coaugmented coassociative coalgebras and their respective $\infty$-categories of modules and comodules following [39, 5.2.], [8]. The next two theorems are [8, Theorem 3.26.]:
@@
(3.1) ${{\mathrm{LMod}}({\mathcal{D}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathrm{Barc}}_{\mathcal{D}}}$${\mathrm{coLMod}({\mathcal{D}}):={\mathrm{LMod}}({\mathcal{D}}^{\mathrm{op}})^{\mathrm{op}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathrm{Barc}}_{\mathcal{C}}}$${\mathrm{co}\mathrm{Alg}({\mathcal{C}})_{{\mathbb{1}}/}:=(\mathrm{Alg}({\mathcal{C}}^{\mathrm{op}})_{/{\mathbb{1}}})^{\mathrm{op}}.}$
@@
Square (3.1) induces on the fiber over every ${\mathrm{A}}\in\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}$ a functor
@@
${\mathrm{Barc}}_{\mathcal{C}}:\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\rightleftarrows\mathrm{co}\mathrm{Alg}({\mathcal{C}})_{{\mathbb{1}}/}:{\mathrm{Cobar}}_{\mathcal{C}}:={\mathrm{Barc}}^{\mathrm{op}}_{{\mathcal{C}}^{\mathrm{op}}},$
${\mathrm{Barc}}_{\mathcal{D}}:{\mathrm{BMod}}({\mathcal{D}})\rightleftarrows\mathrm{coBMod}({\mathcal{D}}):{\mathrm{Cobar}}_{\mathcal{D}}:={\mathrm{Barc}}^{\mathrm{op}}_{{\mathcal{D}}^{\mathrm{op}}}$
and square (3.1) is a map of adjunctions and so induces for any augmented algebra $\mathrm{A}$} in $\mathcal{C}$} and map ${\mathrm{Barc}}_{\mathcal{C}}({\mathrm{A}})\to{\mathrm{A}}^{\prime}$ of coaugmented coalgebras in $\mathcal{C}$} an adjunction
@@
${\mathbb{1}}\otimes_{\mathrm{A}}\mathrm{triv}_{{\mathrm{A}}}({\mathrm{Y}})\simeq\mathrm{colim}_{{n}\in\Delta^{\mathrm{op}}}({\mathrm{A}}^{\otimes{n}}\otimes{\mathrm{Y}})\to{\mathrm{Barc}}_{\mathcal{C}}({\mathrm{A}})\otimes{\mathrm{Y}}\simeq\mathrm{colim}_{{n}\in\Delta^{\mathrm{op}}}({\mathrm{A}}^{\otimes{n}})\otimes{\mathrm{Y}}$
@@
    $(-)^{\vee}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{/{\mathrm{triv}}}\simeq\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})_{/{\mathrm{triv}}}\rightleftarrows\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})_{{\mathrm{triv}}/}\simeq\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}_{{\mathrm{triv}}/}:(-)^{\vee}.$
@@
    There is an adjunction ${\mathrm{LMod}}({\mathcal{C}})\rightleftarrows\mathrm{co}{\mathrm{LMod}}({\mathcal{C}})$ and a map of adjunctions
    ${{\mathrm{LMod}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{co}{\mathrm{LMod}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{/{\mathrm{triv}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}_{{\mathrm{triv}}/}.}$
@@
    $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})={\mathrm{LMod}}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})=\mathrm{co}{\mathrm{LMod}}_{{\mathcal{Q}}}({\mathcal{C}}),$
    where the left adjoint lifts the functor ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ left adjoint to the trivial $\mathcal{O}$}-algebra functor and the right adjoint lifts the functor ${\mathrm{triv}}\circ^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the trivial $\mathcal{Q}$}-coalgebra functor.
@@
    $(-)^{\vee}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{/{\mathrm{triv}}}\rightleftarrows\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}_{{\mathrm{triv}}/}:(-)^{\vee}.$
@@
    ${{\mathrm{RMod}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{coRMod}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{/{\mathrm{triv}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}_{{\mathrm{triv}}/}.}$
@@
    ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\rightleftarrows\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}),$
    where the left adjoint lifts the functor $(-)\circ_{\mathcal{O}}{\mathrm{triv}}:{\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\to{\mathcal{C}}$ left adjoint to the trivial right $\mathcal{O}$}-module functor and the right adjoint lifts the functor $(-)\circ^{{\mathcal{Q}}}{\mathrm{triv}}:\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\to{\mathcal{C}}$ right adjoint to the trivial right $\mathcal{Q}$}-comodule.
@@
    ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}}):{\mathrm{triv}}*^{{\mathcal{Q}}}(-),$
    where the left adjoint lifts the functor ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ left adjoint to the trivial $\mathcal{O}$}-algebra functor and the right adjoint lifts the functor ${\mathrm{triv}}*^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the trivial $\mathcal{Q}$}-coalgebra functor.
@@
${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}}):{\mathrm{triv}}\circ^{{\mathcal{Q}}}(-),$
where the left adjoint lifts the functor ${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ left adjoint to the trivial $\mathcal{O}$}-algebra functor and the right adjoint lifts the functor ${\mathrm{triv}}\circ^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the trivial $\mathcal{Q}$}-coalgebra functor. The forgetful functor $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})$ admits a right adjoint $\mathrm{R}$}. The functor
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ_{\mathcal{O}}(-)}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})\xrightarrow{{\mathrm{R}}}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ^{\mathcal{Q}}(-)}\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}).$
The latter functor lifts the functor $\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to the composition
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}^{\mathrm{dp},{\mathrm{conil}}}_{\mathrm{Q}}}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}}),$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$
lifts the functor ${\mathrm{triv}}*^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ right adjoint to ${\mathrm{triv}}_{\mathcal{Q}}.$
@@
$\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})\subset\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}^{\prime})\subset\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}^{\prime\prime}).$
@@
$\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})\subset\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}^{\prime})$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}^{\prime\prime})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}^{\prime\prime})$
@@
${\mathrm{triv}}*^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$
@@
${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}^{\prime\prime})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}}^{\prime\prime})$
@@
${\mathrm{triv}}\circ_{\mathcal{O}}(-):\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}^{\prime})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}}^{\prime})$
@@
$\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\subset\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}^{\prime})\xrightarrow{{\mathrm{triv}}\circ_{{\mathcal{O}}}(-)}\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}}^{\prime})\xrightarrow{{\mathrm{L}}^{\prime\prime\prime}}\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})$
is left adjoint to the functor ${\mathrm{triv}}\ast^{{\mathcal{Q}}}(-):\mathrm{co}\mathrm{Alg}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}}).$
@@
(3.2) $(-)\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{triv}}:{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\rightleftarrows\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}):(-)\circ^{{\mathcal{Q}}}{\mathrm{triv}}$